# Universally Optimal Periodic Configurations in the Plane

### v3

@@ C.3.1 Satisfying Necessary Conditions
This final inequality is shown by checking that $-2+1123/100e^{-4d/3}-(2429e^{-d})/200+2e^{2d/3}$ is positive with positive derivative at $d=1$ and then applying Lemma 40 to show its derivative is nonnegative for all $d>1$.
@@
follows from checking positivity of $2002+(7397e^{-7d/3})/1800+1621/200e^{-4d/3}-(2429e^{-d})/$ 200 and its derivative at $d=1$, and then applying Lemma 40 to its derivative, $e^{-d}(2429/200-(51779e^{-4d/3})/5400-(1621e^{-d/3})/150)$, as we did with the bound of $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(-1,-1)$. ∎
@@ C.3.2 Computations for the Linear Approximation Bound
by applying Lemma 40 to $-688/45e^{-19d/3}-9377/225e^{-16d/3}-24e^{-3d}+16e^{-7d/3}$. Finally, with the additional assumption that $B<0$, we must show $A^{2}-BC<0$. To bound $B^{2}$ above, we bound $B$ below (since $B<0$). We have
@@
where we obtain the final inequality by checking positivity of $-520+(400652e^{-d})/225-(114472e^{-d/3})/75+520e^{2d/3}$ and its derivative at $d=1$, and applying Lemma 40 to its derivative. Likewise, the derivative of $2A_{l}/B_{l}$ depends only on the sign of
@@ D.1 Basic Lemmas and Other Estimates
The second upper bound follows in a similar fashion using the fact that for $n\geq 1$, the $n$th term is at least as large as the $-(n+1)$th term. ∎
@@
We get an easy lower bound by just taking the $n=0,-1$ terms. For an upper bound, we bound the tail:
@@
$\left\lvert 2\sum_{n\geq 1}\left[1-2an^{2}\right]e^{-an^{2}}\right\rvert$ $\leq 4\sum_{n\geq 1}an^{2}e^{-an^{2}}\leq 4\sum_{n\geq 1}e^{-2an^{2}/3}\leq 4\sum_{n\geq 1}e^{-2an/3}\leq 4\sum_{n\geq 1}e^{-2(9.6)n/3}$ $=\epsilon_{2}$ (144)
@@ D.2.2 Proof of Lemma 18
we apply the bounds from lemmas 139, 41, 140, we obtain
@@
$=\frac{e^{-a/3}}{\pi^{2}}\left[(a-2+\epsilon)(1+\epsilon)-(1-\epsilon_{2})\right].$ (151)
@@
We have from Lemma 41 and 140 that
@@ D.3 Large $a$ and 4 points
Since $u$ decreases as $t_{2}$ increases on $[-1,1]$ with $u(1)=0$, $u(-1)=1/2$, we have $(e^{-ax^{2}}+e^{-a(x-1)^{2}})e^{-3au^{2}}$ increasing in $t_{2}$ and $e^{-a(\frac{1}{2}-x)^{2}}e^{-3a(\frac{1}{2}-u)^{2}}$ decreasing in $t_{2}$. By the same reasoning, $e^{-ax^{2}}e^{-3au^{2}}$ increases in $t_{1}$, while $e^{-a(\frac{1}{2}-x)^{2}}e^{-3a(\frac{1}{2}-u)^{2}}$ and $e^{-a(x-1)^{2}}e^{-3au^{2}}$ are decreasing in $t_{1}$. Finally,
@@
Now take $(x^{2}+3u^{2}-1/4-1/28)e^{-21(x^{2}+3u^{2}-1/4)}$ as a function of $3u^{2}$. Using the fact that $t_{1}<0$ and $t_{2}\geq 1/2$ on the whole rectangle $A$ (so $21/4<x\leq 1/$ 2 and $120\leq 3u^{2}\leq 1/$ 12), we obtain
@@
and so we may apply Lemma 43 to observe $(x^{2}+3u^{2}-1/4-1/28)e^{-21(x^{2}+3u^{2}-1/4)}$ is increasing as a function of $3u^{2}$. Since $3u^{2}$ is decreasing as a function of $t_{2}$, we finally apply the chain rule to see $-(x^{2}+3u^{2}-1/4-1/28)e^{-21(x^{2}+3u^{2}-1/4)}$ is increasing as a function of $t_{2}$ on all of $A$. In the same way, we can check that it is increasing as a function of $t_{1}$, along with the following, analogous claims:
  • The quantity $-\left[(\frac{1}{2}-x)^{2}+3(\frac{1}{2}-u)^{2}-1/4-4/7\right]e^{-21\left[(\frac{1}{2}-x)^{2}+3(\frac{1}{2}-u)^{2}-1/4\right]}$ is decreasing in $t_{1}$ and $t_{2}$.
  • The quantity $-\left[(x-1)^{2}+3u^{2}-1/4-2/5\right]e^{-21\left[(x-1)^{2}+3u^{2}-1/4\right]}$ is decreasing in $t_{1}$ and increasing in $t_{2}$.
@@
[…]
@@
$\frac{\partial\tilde{g}}{\partial t_{1}}\bigg{|}_{P}$$\leq.62b_{1}≤ .$ (167)
$\leq\frac{ae^{-a/4}3(1+\epsilon).62}{\pi^{2}}.$ (168)
@@
$\geq\frac{5*3*39a^{2}e^{-a/4}}{4*41\pi^{2}}(u-(1-u)e^{-3a(1-2u)})$ (172)
$\geq\frac{14*5*3*39a^{2}e^{-a/4}}{100*4*41\pi^{2}}$ (173)
@@
$\frac{\partial^{2}\tilde{F}-\tilde{g}}{\partial t_{1}\partial t_{2}}\bigg{|}_{P}$ $\geq\frac{14*5*3*39a^{2}e^{-a/4}}{100*4*41\pi^{2}}-\frac{ae^{-a/4}3(1+\epsilon)}{\pi^{2}}$ (174)
$=\frac{ae^{-a/4}}{\pi^{2}}\left(\frac{14*5*3*39a}{100*4*41}-3(1+\epsilon)\right).$ (175)
@@ D.4.3 Bounds for proofs of Lemmas 33 and 36
It remains to show that $\tilde{F}_{T}\geq\tilde{g}_{c,d}^{*}$ for various $c,d$ and various line segments in the critical region $[-1,1]\times[-1/2,1/2]$, where $\tilde{g}_{c,d}^{*}$ is an upper bound for $\tilde{g}_{c,d}$ obtained by replacing $a_{i,j}$’s with upper and lower bounds from Lemma 44.
@@
which gives $\tilde{g}_{c,d}(t_{1},t_{2})\leq\tilde{g}_{c,d}^{*}(t_{1},t_{2})$ if $c<t_{1}<0$ and $-1\leq t_{2}<d\leq 0$. To complete the proof of Lemmas 33 and 36 we show inequalities (62) hold with
@@ D.4.4 Computations for proof of Lemma 34
where this last quantity is greater than 00 because it’s true for $a=9.6$ and clearly increasing in $a$.
@@ D.4.5 Positivity of $L_{1}$ for Proposition 35
on $[0,1]\times[-1/2,0]$. We first bound $\frac{\tilde{f_{1}}\tilde{f_{2}}^{\prime}}{\tilde{f_{1}}^{\prime}\tilde{f_{2}}}$below. Using the Lemma 41, $\tilde{f_{1}}\geq e^{-ax^{2}}$. Since $\lvert t_{2}\rvert\leq\frac{1}{2}$, by (137) it follows that
@@
$\frac{\sin(2\pi x)}{x}=\frac{2\pi\sqrt{1-t_{1}^{2}}}{\arccos(t_{1})}$ is concave in $t_{1}$ for $t_{1}\in(-1,1)$. Also, $\frac{u}{\sin(2\pi u)}$ is decreasing in $t_{2}$ for $t_{2}\in(-1,1)$.
@@
Let $\phi(t)=\frac{\sqrt{1-t^{2}}}{\arccos{t_{1}}}$. Then
@@ D.4.6 Log Derivative Estimates for the proof of Lemma 37
Next, we show $L_{2}(t_{1},t_{2}):=\frac{\tilde{f_{1}}(t_{1})}{\tilde{f_{1}}^{\prime}(t_{1})}-t_{1}-\frac{b_{0,0}+a_{0,1}t_{2}+a_{0,2}t_{2}^{2}}{a_{1,0}+a_{0,2}t_{2}}\geq 0$ at $(0,-1/5)$, or equivalently that

### v4

@@ C.3.1 Satisfying Necessary Conditions
This final inequality is shown by checking that $2+1123/100e^{-4d/3}-(2429e^{-d})/200+2e^{2d/3}$ is positive with positive derivative at $d=1$ and then applying Lemma 40 to show its derivative is nonnegative for all $d>1$.
@@
follows from checking positivity of $2+(7397e^{-7d/3})/1800+1621/200e^{-4d/3}-(2429e^{-d})/200$ and its derivative at $d=1$, and then applying Lemma 40 to its derivative, $e^{-d}(2429/200-(51779e^{-4d/3})/5400-(1621e^{-d/3})/150)$, as we did with the bound of $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(-1,-1)$. ∎
@@ C.3.2 Computations for the Linear Approximation Bound
by applying Lemma 40 to $688/45e^{-19d/3}-9377/225e^{-16d/3}-24e^{-3d}+16e^{-7d/3}$. Finally, with the additional assumption that $B<0$, we must show $A^{2}-BC<0$. To bound $B^{2}$ above, we bound $B$ below (since $B<0$). We have
@@
where we obtain the final inequality by checking positivity of $520+(400652e^{-d})/225-(114472e^{-d/3})/75+520e^{2d/3}$ and its derivative at $d=1$, and applying Lemma 40 to its derivative. Likewise, the derivative of $2A_{l}/B_{l}$ depends only on the sign of
@@ D.1 Basic Lemmas and Other Estimates
The second upper bound follows in a similar fashion using the fact that for $n\geq 1$, the $n$th term is at least as large as the (n+1)-(n+1)th term. ∎
@@
We get an easy lower bound by just taking the n=0,1n=0,-1 terms. For an upper bound, we bound the tail:
@@
$\left\lvert 2\sum_{n\geq 1}\left[1-2an^{2}\right]e^{-an^{2}}\right\rvert$ $\leq 4\sum_{n\geq 1}an^{2}e^{-an^{2}}\leq 4\sum_{n\geq 1}e^{-2an^{2}/3}\leq 4\sum_{n\geq 1}e^{-2an/3}\leq 4\sum_{n\geq 1}e^{-2(9.6)n/3}$ $<\epsilon_{2}$ (144)
@@ D.2.2 Proof of Lemma 18
we first use the bounds from inequalities (139), (140) and Lemma 41 to arrive at the following two inequalities:
@@
$=\frac{e^{-a/3}}{\pi^{2}}\left[(a-2+\epsilon)(1+\epsilon)-(1-\epsilon_{2})\right],$ (151)
@@
We have from Lemma 41 and inequality (140) that
@@ D.3 Large $a$ and 4 points
Since $u$ decreases as $t_{2}$ increases on [1,1][-1,1] with $u(1)=0$, u(1)=1/2u(-1)=1/2, we have $(e^{-ax^{2}}+e^{-a(x-1)^{2}})e^{-3au^{2}}$ increasing in $t_{2}$ and $e^{-a(\frac{1}{2}-x)^{2}}e^{-3a(\frac{1}{2}-u)^{2}}$ decreasing in $t_{2}$. By the same reasoning, $e^{-ax^{2}}e^{-3au^{2}}$ increases in $t_{1}$, while $e^{-a(\frac{1}{2}-x)^{2}}e^{-3a(\frac{1}{2}-u)^{2}}$ and $e^{-a(x-1)^{2}}e^{-3au^{2}}$ are decreasing in $t_{1}$. Finally,
@@
Now take $(x^{2}+3u^{2}-1/4-1/28)e^{-21(x^{2}+3u^{2}-1/4)}$ as a function of $3u^{2}$. Using the fact that $t_{1}<0$ and $t_{2}\geq 1/2$ on the whole rectangle $A$ (so $1/4<x\leq 1/2$ and $0\leq 3u^{2}\leq 1/12$), we obtain
@@
and so we may apply Lemma 43 to observe $(x^{2}+3u^{2}-1/4-1/28)e^{-21(x^{2}+3u^{2}-1/4)}$ is increasing as a function of $3u^{2}$. Since $3u^{2}$ is decreasing as a function of $t_{2}$, we finally apply the chain rule to see $(x^{2}+3u^{2}-1/4-1/28)e^{-21(x^{2}+3u^{2}-1/4)}$ is increasing as a function of $t_{2}$ on all of $A$. In the same way, we can check that it is increasing as a function of $t_{1}$, along with the following, analogous claims:
  • The quantity $\left[(\frac{1}{2}-x)^{2}+3(\frac{1}{2}-u)^{2}-1/4-4/7\right]e^{-21\left[(\frac{1}{2}-x)^{2}+3(\frac{1}{2}-u)^{2}-1/4\right]}$ is decreasing in $t_{1}$ and $t_{2}$.
  • The quantity $\left[(x-1)^{2}+3u^{2}-1/4-2/5\right]e^{-21\left[(x-1)^{2}+3u^{2}-1/4\right]}$ is decreasing in $t_{1}$ and increasing in $t_{2}$.
@@
$\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}\bigg|_{P}\geq 0,\hskip 28.45274pt\frac{\partial^{2}\tilde{F}-\tilde{g}}{\partial t_{1}\partial t_{2}}\bigg|_{P}\geq 0.$
@@
$\frac{\partial\tilde{g}}{\partial t_{1}}\bigg|_{P}$ $\leq\frac{62b_{1}}{100}$ (167)
$\leq\frac{ae^{-a/4}3(1+\epsilon)62}{100\pi^{2}}.$ (168)
@@
$\geq\frac{5\times 3\times 39a^{2}e^{-a/4}}{4\times 41\pi^{2}}(u-(1-u)e^{-3a(1-2u)})$ (172)
$\geq\frac{14\times 5\times 3\times 39a^{2}e^{-a/4}}{100\times 4\times 41\pi^{2}}$ (173)
@@
$\frac{\partial^{2}\tilde{F}-\tilde{g}}{\partial t_{1}\partial t_{2}}\bigg|_{P}$ $\geq\frac{14\times 5\times 3\times 39a^{2}e^{-a/4}}{100\times 4\times 41\pi^{2}}-\frac{ae^{-a/4}3(1+\epsilon)}{\pi^{2}}$ (174)
$=\frac{ae^{-a/4}}{\pi^{2}}\left(\frac{14\times 5\times 3\times 39a}{100\times 4\times 41}-3(1+\epsilon)\right).$ (175)
@@ D.4.3 Bounds for proofs of Lemmas 33 and 36
It remains to show that $\tilde{F}_{T}\geq\tilde{g}_{c,d}^{*}$ for various $c,d$ and various line segments in the critical region $1,1]\times[-1/2,1/2]$, where $\tilde{g}_{c,d}^{*}$ is an upper bound for $\tilde{g}_{c,d}$ obtained by replacing $a_{i,j}$’s with upper and lower bounds from Lemma 44.
@@
which gives $\tilde{g}_{c,d}(t_{1},t_{2})\leq\tilde{g}_{c,d}^{*}(t_{1},t_{2})$ if $c<t_{1}<0$ and $1\leq t_{2}<d\leq 0$. To complete the proof of Lemmas 33 and 36 we show inequalities (62) hold with
@@ D.4.4 Computations for proof of Lemma 34
where this last quantity is greater than 0 because it’s true for $a=9.6$ and clearly increasing in $a$.
@@ D.4.5 Positivity of $L_{1}$ for Proposition 35
on $0,1]\times[-1/2,0]$. We first bound $\frac{\tilde{f_{1}}\tilde{f_{2}}^{\prime}}{\tilde{f_{1}}^{\prime}\tilde{f_{2}}}$below. Using the Lemma 41, $\tilde{f_{1}}\geq e^{-ax^{2}}$. Since $\lvert t_{2}\rvert\leq\frac{1}{2}$, by (137) it follows that
@@
Recall $x=\frac{\arccos(t_{1})}{2\pi}$ and $u=\frac{\arccos(t_{2})}{2\pi}$. The function $t_{1}\rightarrow\frac{2\pi\sqrt{1-t_{1}^{2}}}{\arccos(t_{1})}=\frac{\sin(2\pi x)}{x}$ is concave for $t_{1}\in(-1,1)$. Also, $t_{2}\rightarrow\frac{u}{\sin(2\pi u)}$ is decreasing for $t_{2}\in(-1,1)$.
@@
Let $\phi(t)=\frac{\sqrt{1-t^{2}}}{\arccos{t}}$. Then
@@ D.4.6 Log Derivative Estimates for the proof of Lemma 37
Next, we show $L_{2}(t_{1},t_{2}):=\frac{\tilde{f_{1}}(t_{1})}{\tilde{f_{1}}^{\prime}(t_{1})}-t_{1}-\frac{b_{0,0}+a_{0,1}t_{2}+a_{0,2}t_{2}^{2}}{a_{1,0}+a_{0,2}t_{2}}\geq 0$ at (0,1/5)(0,-1/5), or equivalently that
@@
Acknowledgments
The authors thank Henry Cohn, Denali Relles, Larry Rolen, Ed Saff, and Yujian Su for helpful discussions at various stages of this project and also thank the anonymous reviewers for their careful reading and suggestions.